The following information applies to two questions below. Suppose that 3 factors affect the demand for the Bentley Continental in Qatar: Price, income levels (measured by GDP/capita) and number of people living in Qatar. Assume, however, that the number of people living in Qatar, has remained roughly constant during the period 2017-2019.

a. Suppose that in 2017 the price of the Bentley Continental was USD 320,000 in 2017 and that the company sold 25 of these vehicles. For 2018, it kept the price at USD 320,000 while income levels increased by 5%. The company sold 30 Bentley Continentals in 2018. What is your estimate of the income elasticity of demand?

b. In addition to ALL of the information above (including the info in the above question), suppose now that Bentley raised the price of the model to USD 340,000 in 2019, while income levels increased by yet another 5% in 2019 (you can assume that the income elasticity didn’t change between 2018 and 2019). The number of vehicles sold in 2019 was 32. As the brand manager of the Bentley Continental in Qatar, what would be your best estimate of the price elasticity of demand for the vehicle evaluated at the price of USD 320,000 (i.e., at that price, not over a range)?
Answer #1

a) Given, Price in 2017(P1) = USD 320,000

No. of vehicles sold in 2017 (Q1) = 25

Percentage increase in Income level = 5%

No. of cars sold in 2018 = 30

Income Elasticity = Percentage change in Quantity demanded / Percentage change in Income

Q2 - Q1 / Q1 * 100 / 5

= 30 - 25 / 25 *100 / 5

5/25*100 / 5

= 20 /5

Income Elasticity = 4

b) Income Elasticity = 4

Percentage change in Income = 5%

Percentage change in Quantity = 20%

Q1 = 30

Q2 = 32

P1 = 320000

P2 = 340000

Percentage Chane in Price = 340000 - 320000 / 320000 *100

= 20000 / 320000 *100

= 6.25

As per Law of Demand, there should be actual reduction in the Quantity demanded.

At the same time income is rising, so there is a rise in quantity demanded.

Thus overall Change in Quantity, Q1 = 30

Q2 = 27

Price Elasticity = 27 - 30 / 30 * 320000 / 340000 - 320000

= -3 / 30 * 320000 / 20000

= -1.6